OpenStudy (anonymous):

At Walmart, a Panasonic stereo is marked up 25% based on selling price. What is the corresponding percent markup based on cost? Round to the nearest tenth of a percent.

OpenStudy (comm.dan):

I know it has something to do with multiplying @abb

OpenStudy (comm.dan):

I meant @abbiecal

Directrix (directrix):

I think the attachment will assist our thinking about this problem.

OpenStudy (kropot72):

There is not enough information given to answer this question.

OpenStudy (kropot72):

@Directrix Do you agree?

Directrix (directrix):

I'm thinking that the markdown would be 20%.

Directrix (directrix):

Assume: Item cost $1 stereo is marked up 25% --> $1.25 corresponding percent markup based on cost: (1- 1.25)/(1.25) = - .2 = - 20% 20% mark down is just a conjecture.

OpenStudy (kropot72):

There is no way to reference the level of cost to the selling price. Let the selling price = s The amount of mark up = s * (25/100) = s/4 Let the cost price = c Then the percentage mark up based on cost is\[\frac{\frac{s}{4}}{c}\times 100=\frac{s}{4c}\times 100=\frac{100s}{4c}\]

OpenStudy (kropot72):

@Directrix My bad in previous postings. The problem can be solved with the given information. Let the selling price = S The mark up of 25% of S gives an amount S/4. Therefore the cost price = S - S/4 =(3S)/4 The percentage mark up based on the cost price is \[\frac{\frac{S}{4}}{\frac{3S}{4}}\times 100=\frac{S}{4}\times \frac{4}{3S}\times 100=\frac{100}{3}\%\]

Directrix (directrix):

@kropot72 I agree.

OpenStudy (kropot72):

@Directrix Many thanks for your input :)

Directrix (directrix):

@kropot72 I couldn't resist writing it out using that formula. :)